OpenStudy (anonymous):

Combine like terms to simplify the expression: -7x + 3y - 2 + 6x - 1 - y2

OpenStudy (hba):

\[-7x+3y-2+6x-1-y^2\] Can you combine the like terms its easy.

OpenStudy (anonymous):

no.. lol like 7x & 6x? idk

OpenStudy (hba):

Yeah you can combine -7x+6x

OpenStudy (hba):

Also -2-1

OpenStudy (anonymous):

oh ok .. lol

OpenStudy (hba):

So what will be the answer after you combine those up ::P

OpenStudy (anonymous):

-1 & -3?

OpenStudy (hba):

-1x*

OpenStudy (anonymous):

oh lol yeah i forgot .

OpenStudy (hba):

= -7x + 3y - 2 + 6x - 1 - y^2 = -7x+6x-2-1+3y-y^2 = So what will be the answer ?

OpenStudy (anonymous):

idk i just tried and it made lke no sense -.-

OpenStudy (anonymous):

lol

OpenStudy (hba):

(-7x+6x)-(2+1)+3y-y^2 = You have done it do it :P

OpenStudy (anonymous):

-1x - 3 + 3y - y^2 ?

OpenStudy (hba):

Yo you nailed it right there.

OpenStudy (hba):

That is your answer. Anything else i can do for yuo ?

OpenStudy (anonymous):

but im filling in my study guide and thats not one of the multiple choice answers :(

OpenStudy (hba):

What are the options ?

OpenStudy (anonymous):

A) -2y3 + 13x - 3 B) - 3y3 + 13x2 - 3 C) - y2 + 3y + 42x2 - 3 D) - y2 + 3y + 13x - 3

OpenStudy (hba):

-1x - 3 + 3y - y^2 -y^2+3y-x-3 ^ This is what we get. Is your question correct ?

OpenStudy (anonymous):

OMG no . its not -7x in the beginning its 7x im so dumb

OpenStudy (hba):

So let us start again 7x + 3y - 2 + 6x - 1 - y2 7x+6x-2-1+3y

OpenStudy (hba):

What will the answer be ?

OpenStudy (anonymous):

13x

OpenStudy (hba):

What about the others

OpenStudy (anonymous):

i got it :) thank you . gtg thnk you so much for you help

OpenStudy (hba):

Welcome lol :P Catch up withh you later :)